In a vacuum all colors of light travel at a speed of 3 x 10^8 m/sec. The HeNe laser emits a red color with a wavelength of 633 nm. What is the frequency of this light?
1
Expert's answer
2011-04-12T09:32:00-0400
The frequency of the light of red color would be equal to
f = c/ λ = 3 x 108 / 633 x 10-9 = 4.7 x 1014 Hz.